OpenStudy (kitkat16):

The probability that Evan purchases a comic book from the bookstore (event A) is 0.67, and the probability that Glen purchases a comic book from the bookstore (event B) is 0.74. The probability that both Evan and Glen purchase a comic book is 0.2442. Given this information, which statement is true?

OpenStudy (kitkat16):

Events A and B are independent because P(A|B) = P(A) Events A and B are dependent because P(A and B) do not egual P(A) x P(B) Events A and B are dependent because P(A|B)=P(A) I want to say they are dependent because of the last statement "The probability that both Evan and Glen purchase a comic book is 0.2442"

OpenStudy (kitkat16):

Events A and B are dependent because P(A|B)=P(A)

OpenStudy (kitkat16):

Evan and Glen purschasing comic books are independent so Im totally confused.

OpenStudy (kitkat16):

@terenzreignz can u explain this to me? please

terenzreignz (terenzreignz):

hang on...

terenzreignz (terenzreignz):

Okay, two events are indpendent if and only if P(A and B) = P(A)P(B) So, is the probability that they both buy comic books the product of their individual probabilities?

OpenStudy (kitkat16):

yes

OpenStudy (kitkat16):

and that would make them dependent

terenzreignz (terenzreignz):

Yup.

terenzreignz (terenzreignz):

And by which answer choice? :)

terenzreignz (terenzreignz):

Wait, no, the probability that they both buy comic books is short of the what you get when you multiply 0.67 and 0.74

terenzreignz (terenzreignz):

and THAT makes them dependent.

OpenStudy (kitkat16):

Events A and B are dependent because P(A|B)=P(A)

terenzreignz (terenzreignz):

No... I already told you what made them dependent, just put it in symbols ^_^

OpenStudy (kitkat16):

by being short you mean they " do not equal"

terenzreignz (terenzreignz):

yup.

OpenStudy (kitkat16):

0.67*0.74 = 0.4958 0.4958 does not equal 02.442

OpenStudy (kitkat16):

ok THANKS!!! This stuff is confusing. :/

terenzreignz (terenzreignz):

yes

terenzreignz (terenzreignz):

Well, no problem, then ^_^

OpenStudy (kitkat16):

Thanks again!

OpenStudy (anonymous):

Events A and B are dependent because P(A|B)=P(A) is not the answer